$SPAD/src/input lodo2.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool lodo2.output
)set message test on
)set message auto off
)clear all
--S 1 of 26
Q  := Fraction Integer
--R
--R
--R   (1)  Fraction(Integer)
--R                                                                 Type: Domain
--E 1
--S 2 of 26
PQ := UnivariatePolynomial('x, Q)
--R
--R
--R   (2)  UnivariatePolynomial(x,Fraction(Integer))
--R                                                                 Type: Domain
--E 2
--S 3 of 26
x: PQ := 'x
--R
--R
--R   (3)  x
--R                              Type: UnivariatePolynomial(x,Fraction(Integer))
--E 3
--S 4 of 26
Dx: LODO2(Q, PQ) := D()
--R
--R
--R   (4)  D
--IType: LinearOrdinaryDifferentialOperator2(Fraction(Integer),...
--E 4
--S 5 of 26
a := Dx  + 1
--R

--R
--R   (5)  D + 1
--IType: LinearOrdinaryDifferentialOperator2(Fraction(Integer),...
--E 5
--S 6 of 26
b := a + 1/2*Dx**2 - 1/2
--R
--R
--R        1  2       1
--R   (6)  - D  + D + -
--R        2          2
--IType: LinearOrdinaryDifferentialOperator2(Fraction(Integer),...
--E 6
--S 7 of 26
p := 4*x**2 + 2/3
--R
--R
--R          2   2
--R   (7)  4x  + -
--R              3
--R                              Type: UnivariatePolynomial(x,Fraction(Integer))
--E 7
--S 8 of 26
a p
--R
--R
--R          2        2
--R   (8)  4x  + 8x + -
--R                   3
--R                              Type: UnivariatePolynomial(x,Fraction(Integer))
--E 8
--S 9 of 26
(a * b) p = a b p
--R
--R
--R          2         37    2         37
--R   (9)  2x  + 12x + --= 2x  + 12x + --
--R                     3               3
--R                    Type: Equation(UnivariatePolynomial(x,Fraction(Integer)))
--E 9
--S 10 of 26
c := (1/9)*b*(a + b)**2
--R
--R
--R          1  6    5  5   13  4   19  3   79  2    7     1

--R   (10)  -- D  + -- D  + -- D  + -- D  + -- D  + -- D + -
--R         72      36      24      18      72      12     8
--IType: LinearOrdinaryDifferentialOperator2(Fraction(Integer),...
--E 10
--S 11 of 26
(a**2 - 3/4*b + c) (p + 1)
--R
--R
--R           2   44     541
--R   (11)  3x  + -- x + ---
--R                3      36
--R                              Type: UnivariatePolynomial(x,Fraction(Integer))
--E 11
)clear all
--S 12 of 26
PZ   := UnivariatePolynomial(x,Integer)
--R
--R
--R   (1)  UnivariatePolynomial(x,Integer)
--R                                                                 Type: Domain
--E 12
--S 13 of 26
x:PZ := 'x
--R
--R
--R   (2)  x
--R                                        Type: UnivariatePolynomial(x,Integer)
--E 13
--S 14 of 26
Mat  := SquareMatrix(3,PZ)
--R
--R
--R   (3)  SquareMatrix(3,UnivariatePolynomial(x,Integer))
--R                                                                 Type: Domain
--E 14
--S 15 of 26
Vect := DPMM(3, PZ, Mat, PZ);
--R
--R
--R                                                                 Type: Domain
--E 15
--S 16 of 26
Modo := LODO2(Mat, Vect);

--R
--R
--R                                                                 Type: Domain
--E 16
--S 17 of 26
m:Mat := matrix [[x**2,1,0],[1,x**4,0],[0,0,4*x**2]]
--R
--R
--R        + 2         +
--R        |x   1    0 |
--R        |           |
--R   (6)  |     4     |
--R        |1   x    0 |
--R        |           |
--R        |          2|
--R        +0   0   4x +
--R                        Type: SquareMatrix(3,UnivariatePolynomial(x,Integer))
--E 17
--S 18 of 26
p:Vect := directProduct [3*x**2+1,2*x,7*x**3+2*x]
--R
--R
--R           2          3
--R   (7)  [3x  + 1,2x,7x  + 2x]
--IType: DirectProductMatrixModule(3,UnivariatePolynomial(x,Integer),...
--E 18
--S 19 of 26
q: Vect := m * p
--R
--R
--R           4    2        5     2        5     3
--R   (8)  [3x  + x  + 2x,2x  + 3x  + 1,28x  + 8x ]
--IType: DirectProductMatrixModule(3,UnivariatePolynomial(x,Integer),...
--E 19
--S 20 of 26
Dx : Modo := D()
--R
--R
--R   (9)  D
--IType: LinearOrdinaryDifferentialOperator2(SquareMatrix(3,...
--E 20
--S 21 of 26
a : Modo := Dx  + m
--R
--R

--R             + 2         +
--R             |x   1    0 |
--R             |           |
--R   (10)  D + |     4     |
--R             |1   x    0 |
--R             |           |
--R             |          2|
--R             +0   0   4x +
--IType: LinearOrdinaryDifferentialOperator2(SquareMatrix(3,...
--E 21
--S 22 of 26
b : Modo := m*Dx  + 1
--R
--R
--R         + 2         +
--R         |x   1    0 |    +1  0  0+
--R         |           |    |       |
--R   (11)  |     4     |D + |0  1  0|
--R         |1   x    0 |    |       |
--R         |           |    +0  0  1+
--R         |          2|
--R         +0   0   4x +
--IType: LinearOrdinaryDifferentialOperator2(SquareMatrix(3,...
--E 22
--S 23 of 26
c := a*b
--R
--R
--R   (12)
--R   + 2         +     + 4              4    2                  +    + 2         +
--R   |x   1    0 |     |x  + 2x + 2    x  + x            0      |    |x   1    0 |
--R   |           | 2   |                                        |    |           |
--R   |     4     |D  + |   4    2     8     3                   |D + |     4     |
--R   |1   x    0 |     |  x  + x     x  + 4x  + 2        0      |    |1   x    0 |
--R   |           |     |                                        |    |           |
--R   |          2|     |                              4         |    |          2|
--R   +0   0   4x +     +     0            0        16x  + 8x + 1+    +0   0   4x +
--IType: LinearOrdinaryDifferentialOperator2(SquareMatrix(3,...
--E 23
--S 24 of 26
a p
--R
--R
--R            4    2        5     2        5     3      2
--R   (13)  [3x  + x  + 8x,2x  + 3x  + 3,28x  + 8x  + 21x  + 2]
--IType: DirectProductMatrixModule(3,UnivariatePolynomial(x,Integer),...
--E 24

--S 25 of 26
b p
--R
--R
--R            3     2       4         4     3     2
--R   (14)  [6x  + 3x  + 3,2x  + 8x,84x  + 7x  + 8x  + 2x]
--IType: DirectProductMatrixModule(3,UnivariatePolynomial(x,Integer),...
--E 25
--S 26 of 26
(a + b + c) (p + q)
--R
--R
--R   (15)
--R       8      7      6      5      4      3      2
--R   [10x  + 12x  + 16x  + 30x  + 85x  + 94x  + 40x  + 40x + 17,
--R       12      9      8      7     6      5      4      3      2
--R    10x   + 10x  + 12x  + 92x  + 6x  + 32x  + 72x  + 28x  + 49x  + 32x + 19,
--R         8       7        6        5       4       3      2
--R    2240x  + 224x  + 1280x  + 3508x  + 492x  + 751x  + 98x  + 18x + 4]
--IType: DirectProductMatrixModule(3,UnivariatePolynomial(x,Integer),...
--E 26
)spool
)lisp (bye)

———-

References^{[1] nothing}

9